40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 20 23:26:10 2016, gautam, rana, Update, LSC, PSL and AUX-X temperatures changed 2016_01_20_ALS_OutOfLoop_1.pdf
    Reply  Thu Jan 21 18:34:04 2016, rana, Update, LSC, PSL and AUX-X temperatures changed 
       Reply  Fri Jan 22 01:56:01 2016, ericq, Update, LSC, Audio ALSX ALSX.wav
Message ID: 11940     Entry time: Wed Jan 20 23:26:10 2016     Reply to this: 11942  
Author: gautam, rana 
Type: Update 
Category: LSC 
Subject: PSL and AUX-X temperatures changed 

Earlier today, we did a bunch of stuff to see if we could improve the situation with the excess ALS-X noise. Long story short, here are the parameters that were changed, and their initial and final values:

X-end laser diode temperature:     28.5 degrees ---> 31.3 degrees

X-end laser diode current:             1.900 A ---> 1.942 A

X-end laser crystal temperature:   47.43 degrees ---> 42.6 degrees

PSL crystal temperature:              33.43 degrees ---> 29.41 degrees

PSL Diode A temperature:           21.52 degrees ---> 20.75 degrees

PSL Diode B temperature:           22.04 degrees ---> 21.3 degrees 

The Y-end laser temperature has not yet been adjusted - this will have to be done to find the Y-beatnote.

Unfortunately, this does not seem to have fixed the problem - I was able to find the beatnote, with amplitude on the network analyzer in the control room consistent with what we've been seeing over the last few days, but as is clear from Attachment 1, the problem persists...

 


Details:

  • PSL shutter was closed and FSS servo input was turned off. 
  • As I had mentioned in this elog, the beat can now only be found at 47.41 degress +/- 1 deg, which is a shift of almost 5 degrees from the value set sometime last year, ~42.6 degrees. Rana thought it's not a good idea to keep operating the laser at such a high crystal temperature, so we decided to lower the X-end laser temperature back to 42.6 degrees, and then adjust the PSL temperature appropriately such that we found a beat. The diode temperature was also tweaked (this requires using a small screwdriver to twist the little knob inset to the front panel of the laser controller) - for the end laser, we did not have a dedicated power monitor to optimize the diode temperature by maximizing the current, and so we were just doing this by looking at the beat note amplitude on the network analyzer (which wasn't changing by much). So after playing around a little, Rana decided to leave it at 31.3 degrees.
  • We then went to the PSL table and swept through the temperature till a beat was found. The PMC wouldn't stay locked throughout the sweep, so we first did a coarse scan, and saw weak (due to the PMC being locked to some weird mode) beatnotes at some temperatures. We then went back to 29.41 degrees, and ran the PMC autolocker script to lock the PMC - a nice large beatnote was found. 
  • Finally, Rana tweaked the temperatures of the two diodes on the PSL laser controller - here, the optimization was done more systematically, by looking at the PMC transmitted power on the oscilloscope (and also the MEDM screen) as a function of the diode temperature.
  • I took a quick look at the ALS out of loop noise - and unfortunately, our efforts today does not seem to have noticeably improved anything (although the bump at ~1kHz is no longer there). 

Some details not directly related to this work:​ 

  • There are long cables (routed via cable tray) suitable for RF signals that are running from the vertex to either end-table - these are labelled. We slightly re-routed the one running to the X-end, sending it to the IOO rack via the overhead cable tray so that we could send the beat signal from the frequency counter module to the X-end, where we could look at it using an analyzer while also twiddling laser parameters.
  • A webcam (that also claims to have two-way audio!) has been (re?)installed on the PSL table. The ethernet connection to the webcam currently goes to the network switch on the IOO rack (though it is unlabelled at the moment)
  • The X-end area is due for a clean-up, I will try and do some of this tomorrow. 
Attachment 1: 2016_01_20_ALS_OutOfLoop_1.pdf  39 kB  | Hide | Hide all
2016_01_20_ALS_OutOfLoop_1.pdf
ELOG V3.1.3-